OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4

OpenStudy (anonymous):

frist add to eq: then we have: y=4 so for find the x you must y=4 into the one of the eq then we have: y=2/3*4= 8/3 so x=8/3

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

if you use that method, adding them, you actually get \(2y=4\)

OpenStudy (anonymous):

the answer is an ordered pair

OpenStudy (igreen):

Just plug in 2/3x for y in the second equation.. \(y = - \dfrac{2}{3}x + 4\) \(\dfrac{2}{3}x = -\dfrac{2}{3}x + 4\) Add 2/3 to both sides. \(\dfrac{4}{3}x = 4\) Now multiply 3/4 to both sides, what's 3/4 * 4? @skeleking518

OpenStudy (anonymous):

sorry y=2 then x=4/3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

In order to find the solution to the system of equations you have to work on the left hand side of the equation first by using "coefficient matrix" and the right hand side with the answer values.

OpenStudy (igreen):

Yes, so the x-value of our solution is 3. Now we can plug in x = 3 into any of the two equations to solve for y. \(y = \dfrac{2}{3}x\) \(y = \dfrac{2}{3}(3)\) Can you multiply 2/3 * 3? @skeleking518

OpenStudy (anonymous):

2.6 repeating

OpenStudy (igreen):

No, check again.

OpenStudy (anonymous):

3.6 repeating

OpenStudy (igreen):

No..that's wrong.

OpenStudy (anonymous):

3.66666666667

OpenStudy (igreen):

No, \(\dfrac{2}{3} \times 3 = 2\)

OpenStudy (igreen):

So our solution is (3, 2).

OpenStudy (anonymous):

oh

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

thank you

OpenStudy (igreen):

No problem.